# Estimating Dense-Packed Zone Height in Liquid-Liquid Separation: A Physics-Informed Neural Network Approach

### v1

@@ Abstract
Separating liquid–liquid dispersions in gravity settlers is critical in chemical, pharmaceutical, and recycling processes. The dense-packed zone height is an important performance and safety indicator but it is often expensive and impractical to measure due to optical limitations. We propose to estimate phase heights using only inexpensive volume flow measurements. To this end, a physics-informed neural network (PINN) is first pretrained on synthetic data and physics equations derived from a low-fidelity (approximate) mechanistic model to reduce the need for extensive experimental data. While the mechanistic model is used to generate synthetic training data, only volume balance equations are used in the PINN, since the integration of submodels describing droplet coalescence and sedimentation into the PINN would be computationally prohibitive. The pretrained PINN is then fine-tuned with scarce experimental data to capture the actual dynamics of the separator. We then employ the differentiable PINN as a predictive model in an Extended Kalman Filter inspired state estimation framework, enabling the phase heights to be tracked and updated from flow-rate measurements. We first test the two-stage trained PINN by forward simulation from a known initial state against the mechanistic model and a non-pretrained PINN. We then evaluate phase height estimation performance with the filter, comparing the two-stage trained PINN with a two-stage trained purely data-driven neural network. All model types are trained and evaluated using ensembles to account for model parameter uncertainty. In all evaluations, the two-stage trained PINN yields the most accurate phase-height estimates.
@@ 1 Introduction
Estimating the DPZ height by a suitable liquid-liquid gravity separator (LLS) model has the potential to detect flooding caused by DPZ accumulation. The steady-state LLS model of Henschke.1995 remains the state-of-the-art in describing the DPZ distribution along the LLS but dynamic models are limited and not validated (Kamp.2017). Backi2018AFirst-Principles introduced a lumped zero-dimensional LLS model that has the potential of describing DPZ trajectories, yet experimental validation is missing due to lack of suitable data. Both steady-state and lumped models have shown less than 25% accuracy in predicting flow rates at which flooding occurs (Zhai.2025). These shortcomings highlight the need for alternative approaches, such as data-driven modeling. However, purely data-driven methods typically require extensive amounts of data to generalize reliably (shlezinger2022modelbaseddeeplearning).
Physics-informed neural networks (PINNs) (Raissi2019Physics-informedEquations) combine first-principle and data-driven modeling by embedding known physical relationships within the data-driven training process. Thus, PINNs can generalize from scarce training data while adhering to the governing physics. Our previous work (VELIOGLU2025108899) presented an in-silico proof of concept in which mechanistic model data were used as ground truth to demonstrate that a PINN trained with limited data and incomplete physics can predict DPZ height over short time intervals. However, estimation of DPZ height from experimental settler data for longer trajectories remains unexplored.
@@
Once a PINN model is constructed, an estimation scheme is needed to enable real-time estimation of unmeasured states. To this end, Arnold2021StatespaceNetworks showed that PINNs can be effectively integrated within an Extended Kalman Filter (EKF) (Gelb:1974), since Jacobians can be computed through automatic differentiation (naumannAD). Cuomo2024 combined a PINN with the EKF to predict structural displacements in railway motion. In vehicle applications, Tan2023 combined a PINN with the Unscented Kalman Filter (UKF) (JulierUKF) to estimate vehicle attitude, velocity and position with increased accuracy, while decurto2024hybrid demonstrated the use of a PINN-based predictive model as state-transition model in UKF for dynamic models such as a double pendulum system. Fang2025 fine-tuned a pretrained dynamic model with physics guidance to represent hybrid dynamics before embedding them in an EKF framework for vehicle state estimation. These examples suggest that PINNs can serve as surrogates for predictive models when governing equations are incomplete, while the filtering step accounts for measurement noise and model uncertainty. To complement this, ensemble learning (Breiman1996StackedRegressions; dietterich2000ensemble) constitutes an additional technique for improving prediction robustness and uncertainty quantification.
@@
Although PINNs have been combined with Kalman-type filters for state estimation, their application for estimating DPZ heights in pilot-scale LLS has not yet been investigated. Demonstrating that DPZ height can be estimated from readily available volume flow rate measurements with the aid of a PINN would provide significant benefits for LLS operation.
@@
This work aims to estimate the phase heights in a pilot-scale LLS. To this end, we conducted step-response experiments in a pilot-scale LLS by varying the volume flow rate. These experiments capture the transient behavior of DPZ heights and provide data for training, validation, and testing (Section˜2). The low-fidelity mechanistic LLS model (Section˜3) from our previous work (VELIOGLU2025108899) combines simple volume balance equations with complex droplet coalescence and sedimentation submodels. The model is only approximate, as it assumes constant phase heights along the separator length (band-shaped DPZ), neglects axial (convective) transport of the DPZ, droplet–droplet coalescence, and axial and vertical variations in Sauter mean diameter. We generate synthetic pretraining data using the mechanistic model, but use only the volume balance equations as physics regularization in the PINN, as the droplet coalescence and sedimentation submodels are computationally challenging to implement in the PINN due to the necessary discretization of the axial length and droplet classes. A PINN model of the LLS dynamics is then developed by pretraining on synthetic data from the low-fidelity mechanistic model and then fine-tuning with experimental data (Section˜4), similar to the training strategy from Chakraborty2021. The trained PINN is subsequently integrated into an EKF-inspired state estimator to estimate phase heights from volume flow rate measurements (Section˜5). Conclusions and an outlook on future work are presented in Section˜6.
@@ 2.2 Optical measurement of the dense-packed zone height
The phase heights for both the heavy (water) phase and the DPZ are determined from two camera recordings (cf. QIR03, QIR04 in Fig.˜1) taken at four distinct axial positions from each camera. The exact detection positions are shown in Fig.˜2. Similar to the detection algorithm for detecting drop size distributions in endoscope images (Sibirtsev.2023; Sibirtsev.2024) or external camera images (Palmtag.2025), heights are detected by the convolutional neural network based object detection algorithm YOLOv8s trained on manually labelled lab data (Redmon.2016). The manual labelling itself contains errors due to the uneven shape of the drops in the DPZ. Moreover, detection errors can occur due to illumination conditions, often appearing as spikes in the recorded trajectories (cf. Fig.˜3). Furthermore, occasional missing measurements, resulting from object detection errors, appear as discontinuities in the recorded trajectories (cf. Fig.˜3). An exemplary detection is shown in Fig.˜2.
@@
Figure 3: Pre-processed inlet volume flow rate and phase height measurements for the training trajectory, measured at eight different positions at QIR03 and QIR04 (see Fig.˜2).
@@ 2.3 Design of experiments for separator dynamics
Trajectory Time between Volume flow rate $Q_{\textrm{in}}$/ ${\mathrm{m}}^{3}\text{/}\mathrm{h}$
@@
1 – training 2.00 $\mathrm{min}\text{/}$ 1.00 1.50 2.00 1.50 1.00
2 – validation 1.50 $\mathrm{min}\text{/}$ 1.00 1.50 2.00 1.50 1.00
3 – test 2.00 $\mathrm{min}\text{/}$ 1.00 1.25 1.50 1.75 2.00 1.75 1.50 1.25 1.00
4 – test 2.00 $\mathrm{min}\text{/}$ 0.75 1.25 1.75 2.25 1.75 1.25 0.75
@@ 2.4 Post-processing of phase height measurements
We acquired four trajectories with varying time duration from the experimental gravity settler setup, as demonstrated in Section˜2.3. From those, we use one trajectory for training, one trajectory for validation, and two trajectories for testing. All four post-processed trajectories are shown in Fig.˜4. The validation trajectory is used in hyperparameter tuning and tuning for the EKF application (cf. Section˜5.1). The two testing trajectories are acquired to analyze interpolation and extrapolation capabilities of the PINN-based dynamic model. For that reason, we generated training, validation, and interpolation test trajectories with identical bounds for the volume flow rates. However, the extrapolation test trajectory has control actions (and state values) that attain values beyond those bounds. The respective bounds are given in Tab.˜3. The resulting experimental dataset is denoted by $\mathcal{D}_{\mathrm{exp}}$.
@@
Figure 4: Post-processed experimental trajectories obtained from the gravity settler. The quantities $h_{\textrm{HP}}$ and $h_{\textrm{DP}}$ denote the average heavy-phase height and DPZ height, respectively, computed over the eight detection positions. $Q_{\textrm{in}}$ denotes the inlet volume flow rate, and $Q_{\textrm{bot}}$ and $Q_{\textrm{top}}$ denote the bottom and top outlet volume flow rates, respectively.
@@ 3 Low-fidelity mechanistic LLS model
The DPZ is assumed to be band-shaped, meaning constant phase heights, $h_{\textrm{HP}}$ (heavy phase) and $h_{\textrm{DP}}$ (DPZ), along the separator length. Within the DPZ, droplet-droplet coalescence is assumed to be negligible, while droplet-interface coalescence results in a volume flow $Q_{\mathrm{c}}(t)$ into the light phase. The dispersed droplets in the heavy phase rise as a droplet swarm with swarm exponent $n_{\mathrm{swarm}}=2$ (Mersmann.1980; Kampwerth.2020; Ye.2023b), resulting in a volume flow $Q_{\mathrm{s}}(t)$ into the DPZ. Both the coalescence $Q_{\mathrm{c}}(t)$ and sedimentation $Q_{\mathrm{s}}(t)$ flow rates are derived from complex submodels that discretize the axial length of the separator (200) and the droplet classes by diamater (50). Additionally, the volume flow of water from the heavy phase into the DPZ is expressed by the coalescence rate $Q_{\mathrm{c}}(t)$ and the sedimentation rate $Q_{\mathrm{s}}(t)$, by assuming a constant DPZ hold-up of $\bar{\epsilon}_{\mathrm{DP}}=0.9$ in accordance with the separator model proposed by Henschke.1995. The convective transport of the DPZ is also neglected, as well as axial variation of the Sauter mean diameter and its effect on the DPZ.
The separated light and heavy phases exit with outlet volume flow rates $Q_{\textrm{top}}$ and $Q_{\textrm{bot}}$, respectively, both assumed to be free of entrained droplets. The total volume of the three phases is kept constant and equal to the separator volume, yielding the volume balance given by Eq.˜1a. The mechanistic LLS model takes as inputs the initial phase heights $h_{\textrm{HP}}(t=0)$ and $h_{\textrm{DP}}(t=0)$, the inlet volume flow rate $Q_{\textrm{in}}$, the bottom volume flow rate $Q_{\textrm{bot}}$, the Sauter mean diameter at inlet $d_{\mathrm{{32,in}}}$, and provides as outputs the phase heights $h_{\textrm{HP}}(t)$, $h_{\textrm{DP}}(t)$, the bottom outlet volume flow rate $Q_{\textrm{top}}(t)$, the coalescence rate $Q_{\mathrm{c}}(t)$ and sedimentation rate $Q_{\mathrm{s}}(t)$. Further details of the mechanistic LLS model can be found in VELIOGLU2025108899.
@@
Using the mechanistic LLS model described above, we generate 1000 segments with a duration of $1\text{\,}\mathrm{s}\text{/}$ where we sample the model inputs with Latin Hypercube Sampling (LHS) (Iman1981AnAssessment) from the bounds provided in Tab.˜3. We use all of the segments as pretraining data. The simulation segments have a constant step size of $\Delta t=$0.1\text{\,}\mathrm{s}\text{/}$\$ for the PINN (cf. Section˜4). The pretraining dataset is referred to as $\mathcal{D}_{\mathrm{sim}}$. We choose the bounds for initial states and control inputs that correspond to minima and maxima of state values in the experimental trajectories (cf. Tab.˜3), including the extrapolation trajectory.

### v2

@@ Abstract
Separating liquid–liquid dispersions in gravity settlers is critical in chemical, pharmaceutical, and recycling processes. The dense-packed zone height is an important performance and safety indicator but it is often expensive and impractical to measure due to optical limitations. We propose a framework to estimate phase heights by combining a PINN model with readily available volume flow measurements, without requiring phase height measurements during deployment. To this end, a physics-informed neural network (PINN) is first pretrained on synthetic data and physics equations derived from a low-fidelity (approximate) mechanistic model to reduce the need for extensive experimental data. While the mechanistic model is used to generate synthetic training data, only volume balance equations are used in the PINN, as incorporating droplet coalescence and sedimentation submodels would be computationally prohibitive. The pretrained PINN is then fine-tuned with scarce experimental phase height and flow-rate data to capture the actual dynamics of the separator. We then deploy the differentiable PINN as a predictive model in an Extended Kalman Filter inspired state estimation framework, enabling the phase heights to be tracked and updated using flow-rate measurements only. We first test the two-stage trained PINN by forward simulation from a known initial state against the mechanistic model and a non-pretrained PINN. We then evaluate phase height estimation performance with the filter, comparing the two-stage trained PINN with a two-stage trained purely data-driven neural network. All model types are trained and evaluated using ensembles to account for model parameter uncertainty. In all evaluations, the two-stage trained PINN yields the most accurate phase-height estimates.
@@ 1 Introduction
Estimating the DPZ height by a suitable liquid-liquid gravity separator (LLS) model has the potential to detect flooding caused by DPZ accumulation. The steady-state LLS model of Henschke.1995 remains the state-of-the-art in describing the DPZ distribution along the LLS but dynamic models are limited and not validated (Kamp.2017). Backi2018AFirst-Principles introduced a lumped zero-dimensional LLS model that has the potential of describing DPZ trajectories (i.e., temporal evolution of DPZ under dynamic operating conditions), yet experimental validation is missing due to lack of suitable data. Both steady-state and lumped models have shown less than 25% accuracy in predicting flow rates at which flooding occurs (Zhai.2025). These shortcomings highlight the need for alternative approaches, such as data-driven modeling. However, purely data-driven methods typically require extensive amounts of data to generalize reliably (shlezinger2022modelbaseddeeplearning).
Physics-informed neural networks (PINNs) (Raissi2019Physics-informedEquations) combine first-principle and data-driven modeling by embedding known physical relationships within the data-driven training process. Thus, PINNs can generalize from scarce training data while adhering to the governing physics. Our previous work (VELIOGLU2025108899) presented an in-silico proof of concept in which mechanistic model data were used as ground truth to demonstrate that a PINN trained with limited data and incomplete physics can predict DPZ height over short time intervals. However, estimation of DPZ height from experimental settler data for longer time intervals remains unexplored.
@@
While the PINN provides a predictive model of the system dynamics, it cannot by itself reconstruct the true system state under realistic operating conditions, where phase heights are not measurable, initial conditions are unknown, and flow-rate measurements are noisy. Therefore, an estimation scheme is required to fuse model predictions with measurements and enable real-time phase height estimation. To this end, Arnold2021StatespaceNetworks showed that PINNs can be effectively integrated within an Extended Kalman Filter (EKF) (Gelb:1974), since Jacobians can be computed through automatic differentiation (naumannAD). Cuomo2024 combined a PINN with the EKF to predict structural displacements in railway motion. In vehicle applications, Tan2023 combined a PINN with the Unscented Kalman Filter (UKF) (JulierUKF) to estimate vehicle attitude, velocity and position with increased accuracy, while decurto2024hybrid demonstrated the use of a PINN-based predictive model as state-transition model in UKF for dynamic models such as a double pendulum system. Fang2025 fine-tuned a pretrained dynamic model with physics guidance to represent hybrid dynamics before embedding them in an EKF framework for vehicle state estimation. These examples suggest that PINNs can serve as surrogates for predictive models when governing equations are incomplete, while the filtering step accounts for measurement noise and model uncertainty. To complement this, ensemble learning (Breiman1996StackedRegressions; dietterich2000ensemble) constitutes an additional technique for improving prediction robustness and uncertainty quantification.
@@
Although PINNs have been combined with Kalman-type filters for state estimation, their application for estimating DPZ heights in pilot-scale LLS has not yet been investigated. Demonstrating that DPZ height can be estimated from readily available volume flow-rate measurements with the aid of a PINN would provide significant benefits for LLS operation.
@@
This work aims to estimate the phase heights in a pilot-scale LLS. The considered system is a horizontal cylindrical gravity separator (length = 1 m, radius = 0.1 m), where the dispersion flows along the axial direction and separates into a light phase, a dispersion layer (DPZ), and a heavy phase. We conducted step-response experiments in the pilot-scale LLS by varying the volume flow rate. These experiments capture the transient behavior of DPZ heights and provide data for training, validation, and testing (Section˜2). The low-fidelity mechanistic LLS model (Section˜3) from our previous work (VELIOGLU2025108899) combines simple volume balance equations with complex droplet coalescence and sedimentation submodels. The model is only approximate, as it assumes constant phase heights along the separator length (band-shaped DPZ), neglects axial (convective) transport of the DPZ, droplet–droplet coalescence, and axial and vertical variations in Sauter mean diameter. We generate synthetic pretraining data using the mechanistic model, but use only the volume balance equations as physics regularization in the PINN, as the droplet coalescence and sedimentation submodels are computationally challenging to implement in the PINN due to the necessary discretization of the axial length and droplet classes. A PINN model of the LLS dynamics is then developed by pretraining on synthetic data from the low-fidelity mechanistic model and then fine-tuning with experimental phase-height and flow-rate data (Section˜4), similar to the training strategy from Chakraborty2021. The trained PINN is subsequently integrated into an EKF-inspired state estimator to estimate phase heights solely from volume flow-rate measurements during deployment (Section˜5). Conclusions and an outlook on future work are presented in Section˜6.
@@ 2.2 Optical measurement of the dense-packed zone height
The phase heights for both the heavy (water) phase and the DPZ are determined from two camera recordings (cf. QIR03, QIR04 in Fig.˜1) taken at four distinct axial positions from each camera. The exact detection positions are shown in Fig.˜2. Fig.˜3 shows the optically measured phase heights for the training trajectory, where the quantities of interest are the average heavy-phase and DPZ heights over the available detection positions; these averaged phase heights are the variables estimated in the present study.
Similar to the detection algorithm for detecting drop size distributions in endoscope images (Sibirtsev.2023; Sibirtsev.2024) or external camera images (Palmtag.2025), heights are detected by the convolutional neural network based object detection algorithm YOLOv8 trained on manually labelled lab data (Redmon.2016). The manual labelling itself contains errors due to the uneven shape of the drops in the DPZ. Moreover, detection errors can occur due to illumination conditions, often appearing as spikes in the recorded trajectories (cf. Fig.˜3). Furthermore, occasional missing measurements, resulting from object detection errors, appear as discontinuities in the recorded trajectories (cf. Fig.˜3). An exemplary detection is shown in Fig.˜2.
@@
Figure 3: Pre-processed inlet volume flow rate and phase height measurements for the training trajectory, measured at eight different positions at QIR03 and QIR04 (see Fig.˜2). Vertical dashed lines denote the times of inlet volume flow setpoint changes.
@@ 2.3 Design of experiments for separator dynamics
In total, four distinct trajectories were acquired, with detailed step durations and flow-rate levels summarized in Tab.˜2. Before each experiment, the plant was heated and circulated for at least one hour until homogeneous conditions were established. Each trajectory was measured only once as repeated experiments under identical conditions would have required repeating the time-intensive preconditioning of the pilot-scale setup. This results in a limited and noisy dataset, which is addressed by the proposed framework in Sections˜4 and 5. Specifically, data scarcity is mitigated through physics-based regularization and pre-training on synthetic data (cf. Sections˜4.4 and 4.2), while ensemble learning (cf. Section˜4.6) accounts for model uncertainty, and filtering (cf. Section˜5.1) accounts for both model uncertainty and measurement noise. The use of single-run experimental data therefore not only reflects practical constraints, but it is intended to demonstrate that the proposed modeling approach can achieve generalization under data-lean scenarios.
@@
Trajectory Time between Volume flow rate $Q_{\textrm{in}}$/ $\mathrm{L}\text{/}\mathrm{s}$
@@
1 – training 2.00 $\mathrm{min}\text{/}$ 0.278 0.417 0.556 0.417 0.278
2 – validation 1.50 $\mathrm{min}\text{/}$ 0.278 0.417 0.556 0.417 0.278
3 – test (interpolation) 2.00 $\mathrm{min}\text{/}$ 0.278 0.347 0.417 0.486 0.556 0.486 0.417 0.347 0.278
4 – test (extrapolation) 2.00 $\mathrm{min}\text{/}$ 0.208 0.347 0.486 0.625 0.486 0.347 0.208
@@ 2.4 Post-processing of phase height measurements
We acquired four trajectories with varying time duration from the experimental gravity settler setup, as demonstrated in Tab.˜2. From those, we use trajectory #1 (Fig.˜4(a)) for training, trajectory #2 (Fig.˜4(b)) for validation, and trajectories #3 and #4 for testing. Specifically, trajectory #3 (Fig.˜4(c)) is used for the interpolation test, as it contains inlet volume flow rates within the range $Q_{\textrm{in}}\in[0.278,\,0.556]\,$\mathrm{L}\text{/}\mathrm{s}$\$, but has values at intermediate levels not encountered during training (e.g., $0.347\,$\mathrm{L}\text{/}\mathrm{s}$\$ in Tab.˜2). Trajectory #4 (Fig.˜4(d)) is used for the extrapolation test, as it includes inlet volume flow rates $Q_{\textrm{in}}$ outside the range observed in the training trajectory (e.g., $0.625\,$\mathrm{L}\text{/}\mathrm{s}$\$ in Tab.˜2). The validation trajectory is used in hyperparameter tuning and tuning for the EKF application (cf. Section˜5.1). The respective bounds for all variables are given in Tab.˜3. The resulting experimental dataset is denoted by $\mathcal{D}_{\mathrm{exp}}$.
@@
Figure 4: Post-processed experimental trajectories obtained from the gravity settler. The quantities $h_{\textrm{HP}}$ and $h_{\textrm{DP}}$ denote the average heavy-phase height and DPZ height, respectively, computed over the eight detection positions. $Q_{\textrm{in}}$ denotes the inlet volume flow rate, and $Q_{\textrm{bot}}$ and $Q_{\textrm{top}}$ denote the bottom and top outlet volume flow rates, respectively. Vertical dashed lines denote the times of inlet volume flow setpoint changes.
@@ 3 Low-fidelity mechanistic LLS model
The DPZ is assumed to be band-shaped, meaning constant phase heights, $h_{\textrm{HP}}$ (heavy phase) and $h_{\textrm{DP}}$ (DPZ), along the separator length. Within the DPZ, droplet-droplet coalescence is assumed to be negligible, while droplet-interface coalescence results in a volume flow $Q_{\mathrm{c}}(t)$ into the light phase. The dispersed droplets in the heavy phase rise as a droplet swarm with swarm exponent $n_{\mathrm{swarm}}=2$ (Mersmann.1980; Kampwerth.2020; Ye.2023b), resulting in a volume flow $Q_{\mathrm{s}}(t)$ into the DPZ. Both the coalescence $Q_{\mathrm{c}}(t)$ and sedimentation $Q_{\mathrm{s}}(t)$ flow rates are derived from complex submodels that discretize the axial length of the separator into 200 spatial discretization elements and the droplet size distribution into 50 droplet size classes. Additionally, the volume flow of water from the heavy phase into the DPZ is expressed by the coalescence rate $Q_{\mathrm{c}}(t)$ and the sedimentation rate $Q_{\mathrm{s}}(t)$, by assuming a constant DPZ hold-up of $\bar{\epsilon}_{\mathrm{DP}}=0.9$ in accordance with the separator model proposed by Henschke.1995. The convective transport of the DPZ is also neglected, as well as axial variation of the Sauter mean diameter and its effect on the DPZ.
The separated light and heavy phases exit with outlet volume flow rates $Q_{\textrm{top}}$ and $Q_{\textrm{bot}}$, respectively, both assumed to be free of entrained droplets. The total volume of the three phases is kept constant and equal to the separator volume, yielding the volume balance given by Eq.˜1a. The mechanistic LLS model takes as inputs the initial phase heights $h_{\textrm{HP}}(t=0)$ and $h_{\textrm{DP}}(t=0)$, the inlet volume flow rate $Q_{\textrm{in}}$, the bottom volume flow rate $Q_{\textrm{bot}}$, the Sauter mean diameter at inlet $d_{\mathrm{{32,in}}}$, and provides as outputs the phase heights $h_{\textrm{HP}}(t)$, $h_{\textrm{DP}}(t)$, the top outlet volume flow rate $Q_{\textrm{top}}(t)$, the coalescence rate $Q_{\mathrm{c}}(t)$ and sedimentation rate $Q_{\mathrm{s}}(t)$. Further details of the mechanistic LLS model can be found in VELIOGLU2025108899.
@@
Using the mechanistic LLS model described above, we generate 1000 segments with a duration of $1\text{\,}\mathrm{s}\text{/}$ where we sample the model inputs with Latin Hypercube Sampling (LHS) (Iman1981AnAssessment) from the bounds provided in Tab.˜3. We use all of the segments as pretraining data. The simulation segments have a constant step size of $\Delta t=$0.1\text{\,}\mathrm{s}\text{/}$\$ for the PINN (cf. Section˜4). The pretraining dataset is referred to as $\mathcal{D}_{\mathrm{sim}}$. We choose the bounds for initial states and control inputs that correspond to minima and maxima of state values in the experimental trajectories (cf. Tab.˜3), including the extrapolation trajectory. The predictive capability of the low-fidelity mechanistic model is later evaluated by forward simulation against experimental trajectories, alongside the proposed PINN approach (see Section˜4.6).